OpenStudy (anonymous):

How many units apart are the points (-8, -8) and (4, 8)?

OpenStudy (anonymous):

I think the answer is 28.

OpenStudy (anonymous):

use the distance formula :)

OpenStudy (mathteacher1729):

@tornjeansxo do you have graph paper handy?

OpenStudy (anonymous):

Distance formula \[\sqrt{(8--8)^{2}+(4--8)^{2}}\]

OpenStudy (anonymous):

yes, wouldn't it just be 28? i mean, i counted the squares.

OpenStudy (anonymous):

this si the distance formula: http://t1.gstatic.com/images?q=tbn:ANd9GcTm6Me_U_doWwOIGSKd-IVvVSFQt-LfoGsYH9AaJSxGuIgjzhzxQN8gU5t-zA plug in your values to the equation and just solve :)

OpenStudy (anonymous):

@tornjeansxo thats incorrect

OpenStudy (anonymous):

oh okay.

OpenStudy (anonymous):

solve the equation and tell me what you get

OpenStudy (anonymous):

okay would it be 16 - 12 then..

OpenStudy (anonymous):

8 - -8 + 4 - -8

OpenStudy (anonymous):

\[\sqrt{[8- (-8)]^2+[4-(-8)]^2}\]\[\sqrt{(8 +8)^2+(4+8)^2}\]\[\sqrt{16^2+12^2}\]\[\sqrt{256+144}\]\[\sqrt{400}\]\[d=20\]

OpenStudy (mathteacher1729):

Here is why the distance formula works. If you know how to do the Pythag. theorem, it should make sense. :)

OpenStudy (anonymous):

oh wow i see now. so the answer is 20?

OpenStudy (anonymous):

@mathteacher1729 that can also be done but its an extra step @tornjeansxo yes, when you work out the distance formula that is the answer you should get

OpenStudy (mathteacher1729):

@yummydum I find it is often a worthy extra step because it helps students understand why the formula works and it helps them remember it for future reference.

OpenStudy (anonymous):

thank you all so much! if i tag you in my next question, can you help me out too? :)